OpenStudy (anonymous):

A mixed apple vendor starts out with $28 \frac{2}{5}$ apples, sells $5 \frac{1}{2}$ apples to Bob, and then sells $7 \frac{1}{3}$ apples to Jane. After these two sales, how many apples does the vendor have left (expressed as a mixed number)?

OpenStudy (anonymous):

so you have to first change all the bases of the fraction to the same number 28 2/5 into 28 12/30 5 1/2 into 5 15/30 7 1/3 into 7 10/30 you then add together 7 10/30 and 5 15/30 that becomes 12 25/30 you can then convert both 28 12/30 and 12 25/30 into mixed numbers 28 12/30= 852/30 12 25/30=385/30 then subtract \[852/30 - 385/30=467/30\] your final answer should be 467/30 as a mixed number or 15 and 17/30